You are doing research on planet X. The temperature inside the space station is a carefully controlled 28 ∘C and the pressure is 753 mmHg . Suppose that a balloon, which has a volume of 850. mL inside the space station, is placed into the airlock, and floats out to planet X.

If planet X has an atmospheric pressure of 0.145 atm and the volume of the balloon changes to 3.22 L , what is the temperature, in degrees Celsius, on planet X (n does not change)? Express your answer using three significant figures.
Answer #1

P1V1 / P2V2 = T1 / T2

P1 = 753 mmHg = 0.99079 atm; V1 = 850 mL = 0.850 L; T1 = 28 0C = 28 + 273.15 K = 301.15 K

P2 = 0.145 atm; V2 = 3.22 L; T2 = ?

T2 = T1 x [P2V2 / P1V1] = 301.15 K x [0.145 atm x 3.22 L / [0.99079 atm x 0.850 L]

T2 = 166.957603054 K = 273.15 K - 166.957603054 = 106.192396946 0C

temperature, in degrees Celsius, on planet X = 106 0C

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")